OpenStudy (anonymous):

How do you write 3x+4y=10in function notation?

OpenStudy (anonymous):

Solve for y and then replace 'y' with f(x) 3x+4y=10 4y=10−3x 4y=−3x+10 y=−3x+104 y=−3x4+104 y=−34x+52 f(x)=−34x+52

OpenStudy (anonymous):

Make sense?

OpenStudy (luigi0210):

Uhm @Peasant1 I don't think that's right..

OpenStudy (anonymous):

You sure?

OpenStudy (anonymous):

I'm sorry you're right...

OpenStudy (anonymous):

i messed up.

OpenStudy (dan815):

3x+4y=10 3x+4y-10=f(x,y) done

OpenStudy (dan815):

lol

OpenStudy (luigi0210):

thank you dan :/

OpenStudy (anonymous):

haha

OpenStudy (dan815):

hahaha

OpenStudy (dan815):

they make the questions for you guys so ambiguous

OpenStudy (luigi0210):

Well time to hit the old trail

OpenStudy (dan815):

where is this old trail

OpenStudy (anonymous):